The sum of the squares of three consecutive positive odd integers is 35. Find the integers. Write the equation and then solve
Answer by scott8148(5891)

let x= first, so x+2=middle and x+4=last ... x^2+(x+2)^2+(x+4)^2=35 ... 3x^2+12x+20=35

3x^2+12x-15=0 ... (3x-3)(x+5)=0